What is the ratio [Na+]in/[Na+]out created by depolarization in a nerve cell if the final potential is +50 mV at 37 0C?

Not sure how you go about solving this with what's given.
Answer #1

Equilibrium potential

ENa= 61/1 log [Na+]out/[Na+]in

Given

ENa= +50 mV

50/61= log [Na+]out/[Na+]in

0.82= log [Na+]out/[Na+]in

Antilog 0.82= [Na+]out/[Na+]in

6.6= [Na+]out/[Na+]in

1/6.6=[Na+]in/[Na+]out

[Na+]in/[Na+]out=0.15:1